OpenStudy (anonymous):

An urban vehicle weighing 2500 kg hits a stationary 1100kg vehicle. The vehicles stick together and continue to slide 50m to a stop. coefficient of friction = .7 . How fast is the EUV going before the collision

OpenStudy (anonymous):

\[m_1v_1+m_2v_2=(m_1+m_2)v'\]

OpenStudy (anonymous):

\[m_1v_1=(m_1+m_2)v'\]

OpenStudy (anonymous):

\[v_1=\frac{(m_1+m_2)v'}{m_1}\]

OpenStudy (anonymous):

don't know v' so we must find it

OpenStudy (anonymous):

given 50m \[50=\frac{1}{2}at^2 +v_0t+s_0\] \[50=\frac{1}{2}at^2 +v_0t+s_0\] \[v_f=at+v_0\] since we kno wthat final velocity after collision is 0 (stops) \[0=at+v_0 \] \[-at=v_0\]

OpenStudy (anonymous):

don't have time soooo maybe use the time independent equation

OpenStudy (anonymous):

a= .7*3600*g Vf^2 =0 = Vi^2 - 2ad Vi^2 = 2(.7*3600*g)*50

OpenStudy (anonymous):

where'd your acceleration come from?

OpenStudy (anonymous):

\[0=v_0^2+2as\] \[-2ag=v_0^2\] \[\sqrt{-2ag}=v_0\]

OpenStudy (anonymous):

\[\sqrt{-2as}=v_0\]

OpenStudy (anonymous):

so now we can replace this with v_0 and we can use the other equation to get rid of the time

OpenStudy (anonymous):

none of that is necessary.

OpenStudy (anonymous):

I showed you how to find the velocity immediately after the collision. Use the conservation of momentum equation to find initial velocity.

OpenStudy (anonymous):

no you simply stated that the acceleration is the (friction coefficient) times the mass of the object

OpenStudy (anonymous):

was this found out through a fbd?

OpenStudy (anonymous):

whoops, yeah... a= .7g... shouldn't have mass in there...

OpenStudy (anonymous):

a= .7*g Vf^2 =0 = Vi^2 - 2ad Vi^2 = 2(.7*g)*50

OpenStudy (anonymous):



OpenStudy (anonymous):

mu*N is the force ma = mu*N = mu*mg -> a=mu*g

OpenStudy (anonymous):

\[\sum F_x=ma=-f\] \[F_y=ma=0=N-W=\] \[N=W\] \[N=mg\] \[\mu N=.7*mg\]

OpenStudy (anonymous):

alright i wasn't sure how the 50 meters factored in?

OpenStudy (anonymous):

a= .7*g Vf^2 =0 = Vi^2 - 2ad Vi^2 = 2(.7*g)*50

OpenStudy (anonymous):

yeah i thought it was to find acceleration alright thanks

OpenStudy (anonymous):

sure